OpenStudy (anonymous):

A group of college students are volunteering for Help the Homeless during their spring break. They are putting the finishing touches on a house they built. Working alone, Jillian can paint a certain room in 14 hours. Mike can paint the same room in 4 hours. Write an equation that can be used to find how long it will take them working together to paint the room. Then, find how many hours will it take them to paint the room?

OpenStudy (anonymous):

x/14 +x/4 =1 4x+14x=56 18x=56 x=28/9=3 1/9 hours

OpenStudy (anonymous):

\[(\frac{1}{14}+\frac{1}{4})t=1\]

OpenStudy (anonymous):

\[\frac{9}{28}t=1\] \[t=\frac{28}{9}\]